OpenStudy (anonymous):

If sin Θ = 2 over 3 and tan Θ < 0, what is the value of cos Θ?

OpenStudy (anonymous):

sin^2+cos^2=1 => cos=plus or minus sqrt(1-sin^2). tg=sin/cos. sin>0 tg<0 => cos<0 => cos=-sqrt(1-sin^2)=-sqrt(1-4/9)= -sqrt(5)/3